What is the tangential acceleration of the object if it accelerates uniformly in circular







motion with changes in velocity of 80 m/s an in the total changes of time of 25 seconds

1
Expert's answer
2021-11-04T10:15:22-0400

The tangential acceleration of the object can be found as follows:


"a_t=\dfrac{\Delta v}{\Delta t}=\dfrac{80\ \dfrac{m}{s}}{25\ s}=3.2\ \dfrac{m}{s^2}."